# Asymptotic independence in higher dimensions and its implications on risk management

### v2

@@ A.2 Proofs of Section 3
First, we present some auxiliary results required for the proof of Theorem 3.9. The following lemma is from Hashorva and Hüsler [20, Proposition 2.5 and Corollary 2.7].
@@
has a unique solution $\text{e}^{S}\in\mathbb{R}^{d}$ such that
@@
    Suppose $\Sigma^{-1}\boldsymbol{1}\ngtr\boldsymbol{0}$. Then $I\neq\mathbb{I}_{d}$ and for any set $S\neq\mathbb{I}_{d}$ with $I\subseteq S\subseteq\mathbb{I}_{d}$ the equality $\kappa_{\mathbb{I}_{d}}=\kappa_{S}$ holds. For $S\subseteq\mathbb{I}_{d}$ with $S^{c}\cap I\not=\emptyset$ we have $I=I_{S}$ and the inequality $\kappa_{\mathbb{I}_{d}}>\kappa_{S}$ holds.
@@
We start with some preliminary calculations. Suppose $S\subseteq\mathbb{I}_{d}$ with $S^{c}\cap I\not=\emptyset$. Let $\boldsymbol{e}^{*}:=\boldsymbol{e}^{\mathbb{I}_{d}}$ be the unique solution of the quadratic programming problem $\mathcal{P}_{\Sigma^{-1}}$ such that $\kappa_{\mathbb{I}_{d}}=\boldsymbol{e}^{*\top}\Sigma^{-1}\boldsymbol{e}^{*}$, $\boldsymbol{e}^{*}\geq\boldsymbol{1}$ and $[\Sigma^{-1}\boldsymbol{e}^{*}]_{S^{c}}\not=\boldsymbol{0}_{S^{c}}$ since $[\Sigma^{-1}\boldsymbol{e}^{*}]_{I}>\boldsymbol{0}_{I}$ and $S^{c}\cap I\not=\emptyset$ (cf. Lemma A.1). First, define $\widetilde{\boldsymbol{e}}_{S}:=\boldsymbol{e}_{S^{c}}^{*}+[\Sigma^{-1}]_{S^{c}}^{-1}[\Sigma^{-1}]_{S^{c}S}\boldsymbol{e}_{S}^{*}$ and note that
@@
Finally, the Schur decomposition (see Lauritzen [27, eq. (B2)])
@@
$\kappa_{\mathbb{I}_{d}}=$ $\boldsymbol{e}^{*\top}\Sigma^{-1}\boldsymbol{e}^{*}$
$=$ $\boldsymbol{e}_{S}^{*\top}\Sigma_{S}^{-1}\boldsymbol{e}^{*}_{S}+\widetilde{\boldsymbol{e}}_{S}^{\top}[\Sigma^{-1}]_{S^{c}}\widetilde{\boldsymbol{e}}_{S}$ (A.2)
@@
(a) If $\Sigma^{-1}\boldsymbol{1}>\boldsymbol{0}$ then $I=\mathbb{I}_{d}$ and $\boldsymbol{e}^{*}=\boldsymbol{1}$; see Hashorva and Hüsler [20, Proposition 2.5]. Thus, any $S\subseteq\mathbb{I}_{d}$ with $S\neq\mathbb{I}_{d}$ satisfies $S^{c}\cap I\not=\emptyset$ and the result follows from (A.3).
(b) If $\Sigma^{-1}\boldsymbol{1}\ngtr\boldsymbol{0}$ then $I\subseteq\mathbb{I}_{d}$ and $I\neq\mathbb{I}_{d}$; see Hashorva and Hüsler [20, Proposition 2.5]. Hence, Lemma A.1 and $\Sigma_{I}^{-1}\boldsymbol{1}_{I}>\boldsymbol{0}_{I}$ imply that
@@
$\Leftarrow$:   Suppose $\Sigma^{-1}\boldsymbol{1}\ngeq\boldsymbol{0}$. From Lemma A.2(b) we already know that $I\not=\mathbb{I}_{d}$. Now let $\ell\in\mathbb{I}_{d}\backslash I$. For $S=\mathbb{I}_{d}\backslash\{\ell\}$ we have $I\subseteq S\subseteq\mathbb{I}_{d}$, with $I=I_{S}$ and $\kappa_{\mathbb{I}_{d}}=\kappa_{S}$ (cf. proof of Lemma A.2). Now using (A.4) we have
@@
$\Rightarrow$:   Suppose there exists $\ell\in\mathbb{I}_{d}$ such that (A.6) holds. We prove the statement by contradiction. By way of contradiction assume $\Sigma^{-1}\boldsymbol{1}>\boldsymbol{0}$ holds. Lemma A.2 says that for any set $S\subseteq\mathbb{I}_{d}$ with $S\not=\mathbb{I}_{d}$ the inequality $\kappa_{\mathbb{I}_{d}}>\kappa_{S}$ holds. Again using (A.4) we have with $\kappa^{*}:=\kappa_{\mathbb{I}_{d}}-\kappa_{\mathbb{I}_{d}\setminus\{\ell\}}$ and $d^{*}:=d-|I_{\mathbb{I}_{d}\setminus\{\ell\}}|$,
@@ A.3 Proofs of Section 4
$\Gamma_{\rho}^{-1}=\begin{bmatrix}\boldsymbol{I}_{k}+\frac{\rho^{2}}{1-k\rho^{2}}\boldsymbol{1}_{k}\boldsymbol{1}_{k}^{\top}\;\;\;\;\;&\frac{-\rho}{1-k\rho^{2}}\boldsymbol{1}_{k}\\[10.00002pt] \frac{-\rho}{1-k\rho^{2}}\boldsymbol{1}_{k}^{\top}&\frac{1}{1-k\rho^{2}}\end{bmatrix}.$ (A.7)
@@
  • If $k+1\notin S$, then $[\Gamma_{\rho}]_{S}={\boldsymbol{I}}_{j}$ and hence
    $[\Gamma_{\rho}]_{S}^{-1}\boldsymbol{1}_{j}=\boldsymbol{1}_{j}>\boldsymbol{0}_{j}.$
Thus Theorem 3.9 implies then that $\boldsymbol{X}_{J}$, for any $J\subseteq\{1,\ldots,k+1\}$ with $|J|\leq k$, is a mutually asymptotically independent random vector in $\mathbb{R}^{J}$. Finally, a conclusion of Lemma 4.2 is that $\boldsymbol{X}$ is $k$-wise asymptotically independent in $\mathbb{R}^{(k+1)\times(k+1)}$, although it is not $(k+1)$-wise asymptotically independent. From Lemma A.1 we know that $I_{\{1,\ldots,k+1\}}=\{1,\ldots,k\}=I_{\{1,\ldots,k\}}$, $\kappa_{\{1,\ldots,k+1\}}=\kappa_{\{1,\ldots,k\}}=k$, $h_{i}^{S_{1}}=h_{i}^{S_{2}}=1$ for $i\in\{1,\ldots,k\}$ and finally, from Proposition A.3 that
@@
After all, define the $(d\times d)$-dimensional correlation $\Sigma_{\rho}$ as a block diagonal matrix having in the first $(d-(k+1))\times(d-(k+1))$ block the identity matrix, zeros in the two off-diagonal blocks, and, in the last $(k+1)\times(k+1)$ block $\Gamma_{\rho}$ with $\rho\in(\frac{1}{k},\min(\frac{1}{k-1},\frac{1}{\sqrt{k}}))$, i.e., the random vector $\boldsymbol{Z}^{*}=(Z_{1}^{*},\ldots,Z_{d}^{*})$ with Gaussian copula $C^{\Sigma_{\rho}}$ has the property that $Z_{1}^{*},\ldots,Z_{d-(k+1)}^{*}$ are an independent sequence which is as well independent of the random vector $\boldsymbol{X}=(Z_{d-k}^{*},\ldots,Z_{d}^{*})$ in $\mathbb{R}^{(k+1)\times(k+1)}$ with Gaussian copula $C^{\Gamma_{\rho}}$. Then by analogous arguments as above, $\boldsymbol{Z}^{*}$ is a $k$-wise asymptotic independent random vector in $\mathbb{R}^{d}$ although it is not $(k+1)$-wise asymptotic independent and
@@ A.4 Proofs of Section 5
$=\frac{\widehat{C}_{J^{*}}(\gamma,\ldots,\gamma)}{\widehat{C}_{J}(\gamma,\ldots,\gamma)},$ (A.8)
@@
    If $d=2$, there is nothing else to prove. Hence, now assume $d\geq 3$. Since $0\leq\mathrm{CTP}_{\gamma}(\boldsymbol{Z}_{1|J})\leq 1$, to show (b), it is sufficient to provide an example of $F\in\mathcal{N}_{\ell}\subseteq\mathcal{N}_{k}\subseteq\mathcal{P}_{k}$ for $k\in\{2,\ldots,\ell\}$, such that for $\boldsymbol{Z}\sim F$, we have $\lim_{\gamma\downarrow 0}\mathrm{CTP}_{\gamma}(\boldsymbol{Z}_{1|J})=1$ To this end, we will choose $F$ with Gaussian copula $C^{\Sigma}$ and positive-definite correlation matrix $\Sigma$ as identified using Theorem 4.5, such that $F$ exhibits $\ell$-wise asymptotic independence but not $(\ell+1)$-wise asymptotic independence and for any $x>0$,
@@
    Hence, $F\in\mathcal{N}_{\ell}$ and by (A.8) we have as well
@@
$\sup_{\gamma\in(0,x^{-1}]}\frac{\widehat{C}_{S}^{\Sigma}(x\gamma,\gamma,\ldots,\gamma)}{\widehat{C}_{S}^{\Sigma}(\gamma,\gamma,\ldots,\gamma)}<\infty$ (A.9)
@@
$\text{CoVaR}_{\gamma_{1},\gamma_{2}}(\boldsymbol{Z}_{1|J})$
$\quad=\inf\{z\in\mathbb{R}_{+}:\mathbb{P}(Z_{1}>z|Z_{j}>\mathrm{VaR}_{\gamma_{2}}(Z_{j}),\,\forall j\in J)\leq\gamma_{1}\}$
$\quad=\mathrm{VaR}_{\gamma_{2}}(Z_{1})\inf\{z\in\mathbb{R}_{+}:\mathbb{P}(Z_{1}>z\mathrm{VaR}_{\gamma_{2}}(Z_{1})|Z_{j}>\mathrm{VaR}_{\gamma_{2}}(Z_{j}),\,\forall j\in J)\leq\gamma_{1}\}.$
Suppose $Z_{1}$ is Pareto$(\alpha)$-distributed, $\alpha>0$. Then the previous equation reduces to
$\text{CoVaR}_{\gamma_{1},\gamma_{2}}(\boldsymbol{Z}_{1|J})$ $=$ $\mathrm{VaR}_{\gamma_{2}}(Z_{1})\inf\{z\in\mathbb{R}_{+}:\frac{\widehat{C}_{J^{*}}(z^{-\frac{1}{\alpha}}\gamma_{2},\gamma_{2},\ldots,\gamma_{2})}{\widehat{C}_{J}(\gamma_{2},\ldots,\gamma_{2})}\leq\gamma_{1}\},$ (A.10)
@@
    $K$ $:=\sup_{\gamma\in(0,\epsilon^{1/\alpha}]}\frac{\widehat{C}_{J^{*}}(\epsilon^{-1/\alpha}\gamma,\gamma,\ldots,\gamma)}{\widehat{C}_{J^{*}}(\gamma,\gamma,\ldots,\gamma)},$ (A.11)
@@
    $\frac{\widehat{C}_{J^{*}}(\gamma,\gamma,\ldots,\gamma)}{\widehat{C}_{J}(\gamma,\gamma,\ldots,\gamma)}\leq\frac{\epsilon}{K},\quad\forall\,\gamma\in(0,\gamma_{0}(\epsilon)).$ (A.12)
    Therefore from (A.11) and (A.12), for all $0<\gamma_{2}<\min(\epsilon^{1/\alpha},\gamma_{0}(\epsilon))$ we have
@@
    Since $\epsilon\in(0,\gamma_{1})$ is arbitrary, this results in
    $\lim_{\gamma_{2}\downarrow 0}\frac{\text{CoVaR}_{\gamma_{1},\gamma_{2}}(\boldsymbol{Z}_{1|J})}{\mathrm{VaR}_{\gamma_{2}}(Z_{1})}=0.$
    Finally, from Lemma 5.5 we already know that $\mathcal{N}_{k}^{*}\subseteq\mathcal{P}_{k}^{*}$, thus the result is true for $\mathcal{N}_{k}^{*}$ as well.
@@
    $\lim_{u\downarrow 0}\frac{\widehat{C}^{\Sigma}_{J^{*}}(xu,u,\ldots,u)}{\widehat{C}^{\Sigma}_{J}(u,\ldots,u)}=1.$ (A.13)
    Additionally, suppose that the margin $F_{1}$ is Pareto($\alpha$)-distributed. Then $F\in\mathcal{N}_{\ell}^{*}\subseteq\mathcal{N}_{k}^{*}\subseteq\mathcal{P}_{k}^{*}$ for $k\in\{2,\ldots,\ell\}$. Due to (A.13), for any $M>0$ there exists an $\gamma_{0}(M)\in(0,1)$ such that
@@
    $\frac{\text{CoVaR}_{\gamma_{1},\gamma_{2}}(\boldsymbol{Z}_{1|J})}{\mathrm{VaR}_{\gamma_{2}}(Z_{1})}$ $=$ $\inf\{z\in\mathbb{R}_{+}:\frac{\widehat{C}_{J^{*}}(z^{-\frac{1}{\alpha}}\gamma_{2},\gamma_{2},\ldots,\gamma_{2})}{\widehat{C}_{J}(\gamma_{2},\ldots,\gamma_{2})}\leq\gamma_{1}\}\geq M,$

### v3

@@ Appendix A Proofs of Section 3
First, we present some auxiliary results required for the proof of Theorem 3.9. The following lemma is from Hashorva and Hüsler [21, Proposition 2.5 and Corollary 2.7].
@@
has a unique solution $\boldsymbol{e}^{S}\in\mathbb{R}^{d}$ such that
@@
    Suppose $\Sigma^{-1}\boldsymbol{1}\ngtr\boldsymbol{0}$. Then $I\neq\mathbb{I}_{d}$ and for any set $S\neq\mathbb{I}_{d}$ with $I\subseteq S\subseteq\mathbb{I}_{d}$ the equality $\kappa_{\mathbb{I}_{d}}=\kappa_{S}$ holds. For $S\subseteq\mathbb{I}_{d}$ with $S^{c}\cap I\not$$\emptyset$ we have $I=I_{S}$ and the inequality $\kappa_{\mathbb{I}_{d}}>\kappa_{S}$ holds.
@@
We start with some preliminary calculations. Suppose $S\subseteq\mathbb{I}_{d}$ with $S^{c}\cap I\not$$\emptyset$. Let $\boldsymbol{e}^{*}:=\boldsymbol{e}^{\mathbb{I}_{d}}$ be the unique solution of the quadratic programming problem $\mathcal{P}_{\Sigma^{-1}}$ such that $\kappa_{\mathbb{I}_{d}}=\boldsymbol{e}^{*\top}\Sigma^{-1}\boldsymbol{e}^{*}$, $\boldsymbol{e}^{*}\geq\boldsymbol{1}$ and $\Sigma^{-1}\boldsymbol{e}^{*}]_{S^{c}}\not=\boldsymbol{0}_{S^{c}}$ since $\Sigma^{-1}\boldsymbol{e}^{*}]_{I}>\boldsymbol{0}_{I}$ and $S^{c}\cap I\not$$\emptyset$ (cf. Lemma A.1). First, define $\widetilde{\boldsymbol{e}}_{S}:=\boldsymbol{e}_{S^{c}}^{*}+[\Sigma^{-1}]_{S^{c}}^{-1}[\Sigma^{-1}]_{S^{c}S}\boldsymbol{e}_{S}^{*}$ and note that
@@
Finally, the Schur decomposition (see Lauritzen [28, eq. (B2)])
@@
$\kappa_{\mathbb{I}_{d}}$ $=$ $\boldsymbol{e}^{*\top}\Sigma^{-1}\boldsymbol{e}^{*}$ (A.2)
$=$ $\boldsymbol{e}_{S}^{*\top}\Sigma_{S}^{-1}\boldsymbol{e}^{*}_{S}+\widetilde{\boldsymbol{e}}_{S}^{\top}[\Sigma^{-1}]_{S^{c}}\widetilde{\boldsymbol{e}}_{S}$
@@
(a) If $\Sigma^{-1}\boldsymbol{1}>\boldsymbol{0}$ then $I=\mathbb{I}_{d}$ and $\boldsymbol{e}^{*}=\boldsymbol{1}$; see Hashorva and Hüsler [21, Proposition 2.5]. Thus, any $S\subseteq\mathbb{I}_{d}$ with $S\neq\mathbb{I}_{d}$ satisfies $S^{c}\cap I\not$$\emptyset$ and the result follows from (A.3).
(b) If $\Sigma^{-1}\boldsymbol{1}\ngtr\boldsymbol{0}$ then $I\subseteq\mathbb{I}_{d}$ and $I\neq\mathbb{I}_{d}$; see Hashorva and Hüsler [21, Proposition 2.5]. Hence, Lemma A.1 and $\Sigma_{I}^{-1}\boldsymbol{1}_{I}>\boldsymbol{0}_{I}$ imply that
@@
[…]
@@
$\Rightarrow$:   Suppose there exists $\ell\in\mathbb{I}_{d}$ such that (A.6) holds. We prove the statement by contradiction. By way of contradiction, assume $\Sigma^{-1}\boldsymbol{1}>\boldsymbol{0}$ holds. Lemma A.2 says that for any set $S\subseteq\mathbb{I}_{d}$ with $S\not=\mathbb{I}_{d}$ the inequality $\kappa_{\mathbb{I}_{d}}>\kappa_{S}$ holds. Again using (A.4) we have with $\kappa^{*}:=\kappa_{\mathbb{I}_{d}}-\kappa_{\mathbb{I}_{d}\setminus\{\ell\}}$ and $d^{*}:=d-|I_{\mathbb{I}_{d}\setminus\{\ell\}}|$,
@@ Appendix B Proofs of Section 4
$\Gamma_{\rho}^{-1}=\begin{bmatrix}\boldsymbol{I}_{k}+\frac{\rho^{2}}{1-k\rho^{2}}\boldsymbol{1}_{k}\boldsymbol{1}_{k}^{\top}\;\;\;\;\;&\frac{-\rho}{1-k\rho^{2}}\boldsymbol{1}_{k}\\[10.00002pt] \frac{-\rho}{1-k\rho^{2}}\boldsymbol{1}_{k}^{\top}&\frac{1}{1-k\rho^{2}}\end{bmatrix}$. (B.1)
@@
  • If $k+1\notin S$, then $\Gamma_{\rho}]_{S}={\boldsymbol{I}}_{j}$ and hence
    $\Gamma_{\rho}]_{S}^{-1}\boldsymbol{1}_{j}=\boldsymbol{1}_{j}>\boldsymbol{0}_{j}.$
Thus Theorem 3.9 implies then that $\boldsymbol{X}_{J}$, for any $J\subseteq\{1,\ldots,k+1\}$ with $|J|\leq k$, is a mutually asymptotically independent random vector in $\mathbb{R}^{J}$. Finally, a conclusion of Proposition 4.3 is that $\boldsymbol{X}$ is $k$-wise asymptotically independent in $\mathbb{R}^{(k+1)\times(k+1)}$, although it is not $(k+1)$-wise asymptotically independent. From Lemma A.1 we know that $I_{\{1,\ldots,k+1\}}=\{1,\ldots,k\}=I_{\{1,\ldots,k\}}$, $\kappa_{\{1,\ldots,k+1\}}=\kappa_{\{1,\ldots,k\}}=k$, $h_{i}^{\{1,\ldots,k+1\}}=h_{i}^{\{1,\ldots,k\}}=1$ for $i\in\{1,\ldots,k\}$ and finally, from Proposition A.3 that
@@
After all, define the $(d\times d)$-dimensional correlation $\Sigma_{\rho}$ as a block diagonal matrix having in the first $(d-(k+1))\times(d-(k+1))$ block the identity matrix, zeros in the two off-diagonal blocks, and, in the last $(k+1)\times(k+1)$ block $\Gamma_{\rho}$ with $\rho\in(\frac{1}{k},\min(\frac{1}{k-1},\frac{1}{\sqrt{k}}))$, i.e., the random vector $\boldsymbol{Z}^{*}=(Z_{1}^{*},\ldots,Z_{d}^{*})$ with Gaussian copula $C^{\Sigma_{\rho}}$ has the property that $Z_{1}^{*},\ldots,Z_{d-(k+1)}^{*}$ are an independent sequence which is as well independent of the random vector $\boldsymbol{X}=(Z_{d-k}^{*},\ldots,Z_{d}^{*})$ in $\mathbb{R}^{(k+1)\times(k+1)}$ with Gaussian copula $C^{\Gamma_{\rho}}$. Then by analogous arguments as above, $\boldsymbol{Z}^{*}$ is a $k$-wise asymptotically independent random vector in $\mathbb{R}^{d}$ although it is not $(k+1)$-wise asymptotically independent and
@@ Appendix C Proofs of Section 5
$=\frac{\widehat{C}_{J^{*}}(\gamma,\ldots,\gamma)}{\widehat{C}_{J}(\gamma,\ldots,\gamma)},$ (C.1)
@@
    If $d=2$, there is nothing else to prove. Hence, now assume $d\geq 3$. Since $0\leq\mathrm{CTP}_{\gamma}(\boldsymbol{Z}_{1|J})\leq 1$, to show (b), it is sufficient to provide an example of $F\in\mathcal{N}_{\ell}\subseteq\mathcal{N}_{k}\subseteq\mathcal{P}_{k}$ for $k\in\{2,\ldots,\ell\}$, such that for $\boldsymbol{Z}\sim F$, we have $\lim_{\gamma\downarrow 0}\mathrm{CTP}_{\gamma}(\boldsymbol{Z}_{1|J})=1$ . To this end, we will choose $F$ with a Gaussian copula $C^{\Sigma}$ and positive-definite correlation matrix $\Sigma$ as identified in Theorem 4.4, such that $F$ exhibits $\ell$-wise asymptotic independence but not $(\ell+1)$-wise asymptotic independence and for any $x>0$,
@@
    Hence, $F\in\mathcal{N}_{\ell}$ and by (C.1) we have as well
@@
$\sup_{\gamma\in(0,x^{-1}]}\frac{\widehat{C}_{S}^{\Sigma}(x\gamma,\gamma,\ldots,\gamma)}{\widehat{C}_{S}^{\Sigma}(\gamma,\gamma,\ldots,\gamma)}<\infty$ (C.2)
@@
$\text{CoVaR}_{\gamma_{1},\gamma_{2}}(\boldsymbol{Z}_{1|J})$ $=$ $\inf\{z\in\mathbb{R}_{+}:\mathbb{P}(Z_{1}>z|Z_{j}>\mathrm{VaR}_{\gamma_{2}}(Z_{j}),\,\forall j\in J)\leq\gamma_{1}\}$
$=$ $\mathrm{VaR}_{\gamma_{2}}(Z_{1})\inf\{z\in\mathbb{R}_{+}:\mathbb{P}(Z_{1}>z\mathrm{VaR}_{\gamma_{2}}(Z_{1})|Z_{j}>\mathrm{VaR}_{\gamma_{2}}(Z_{j}),\,\forall j\in J)\leq\gamma_{1}\}.$
[…]
$\text{CoVaR}_{\gamma_{1},\gamma_{2}}(\boldsymbol{Z}_{1|J})=\mathrm{VaR}_{\gamma_{2}}(Z_{1})\inf\left\{z\in\mathbb{R}_{+}:\frac{\widehat{C}_{J^{*}}(z^{-\frac{1}{\alpha}}\gamma_{2},\gamma_{2},\ldots,\gamma_{2})}{\widehat{C}_{J}(\gamma_{2},\ldots,\gamma_{2})}\leq\gamma_{1}\right\},$ (C.3)
@@
    $K$ $:=\sup_{\gamma\in(0,\epsilon^{1/\alpha}]}\frac{\widehat{C}_{J^{*}}(\epsilon^{-1/\alpha}\gamma,\gamma,\ldots,\gamma)}{\widehat{C}_{J^{*}}(\gamma,\gamma,\ldots,\gamma)},$ (C.4)
@@
    $\frac{\widehat{C}_{J^{*}}(\gamma,\gamma,\ldots,\gamma)}{\widehat{C}_{J}(\gamma,\gamma,\ldots,\gamma)}\leq\frac{\epsilon}{K},\quad\forall\,\gamma\in(0,\gamma_{0}(\epsilon)).$ (C.5)
    Therefore from (C.4) and (C.5), for all $0<\gamma_{2}<\min(\epsilon^{1/\alpha},\gamma_{0}(\epsilon))$ we have
@@
    Since $\epsilon\in(0,\gamma_{1})$ is arbitrary, this results in lim_γ_2↓0 CoVaRγ1,γ2(Z1—J)VaRγ2(Z1)=0. Finally, from Lemma 5.3 we already know that $\mathcal{N}_{k}^{*}\subseteq\mathcal{P}_{k}^{*}$, thus the result is true for $\mathcal{N}_{k}^{*}$ as well.
@@
    $\lim_{u\downarrow 0}\frac{\widehat{C}^{\Sigma}_{J^{*}}(xu,u,\ldots,u)}{\widehat{C}^{\Sigma}_{J}(u,\ldots,u)}=1.$ (C.6)
    Additionally, suppose that the margin $F_{1}$ is Pareto($\alpha$)-distributed. Then $F\in\mathcal{N}_{\ell}^{*}\subseteq\mathcal{N}_{k}^{*}\subseteq\mathcal{P}_{k}^{*}$ for $k\in\{2,\ldots,\ell\}$. Due to (C.6), for any $M>0$ there exists an $\gamma_{0}(M)\in(0,1)$ such that
@@
    $\frac{\text{CoVaR}_{\gamma_{1},\gamma_{2}}(\boldsymbol{Z}_{1|J})}{\mathrm{VaR}_{\gamma_{2}}(Z_{1})}$ $=$ $\inf\left\{z\in\mathbb{R}_{+}:\frac{\widehat{C}_{J^{*}}(z^{-\frac{1}{\alpha}}\gamma_{2},\gamma_{2},\ldots,\gamma_{2})}{\widehat{C}_{J}(\gamma_{2},\ldots,\gamma_{2})}\leq\gamma_{1}\right\}\geq M,$